# Multi-Level Monte Carlo Training of Neural Operators

### v1

@@ 2.2 Mini-batch strategies
$\mathbb{E}_{a\sim\mu}[\nabla_{\theta}X_{\theta}^{m}(a)]\approx\frac{1}{B_{1}}\sum_{j=1}^{B_{1}}[\nabla_{\theta}X_{\theta}^{1}(a_{1,j})]+\sum_{i=2}^{m}\left(\frac{1}{B_{i}}\sum_{j=1}^{B_{i}}[\nabla_{\theta}X_{\theta}^{i}(a_{i,j})-\nabla_{\theta}X_{\theta}^{i-1}(a_{i,j})]\right),$ (3)
where $B_{i}$ stands for the batch size at correction level $i$ in the sum hierarchy, which is obtained by drawing samples from a subset of $N_{i}$ observations at resolution $R_{i}$, for $1\leq i\leq m$. The correction term $-\nabla_{\theta}X_{\theta}^{i-1}(a_{i,j})$ in Eq. 3 is computed by coarsening the sample $a_{i,j}$ to resolution $R_{i-1}$ and evaluating the neural operator. Following Section 2.1, we decrease $B_{i}$ and $N_{i}$ with respect to the level $i$ as the data resolution $R_{i}$ increases.
Sample and batch sizes.
@@
  • Geometric progression. There is a constant growing factor $\delta\geq 1$ between a level $i$ and a subsequent level $i+1$ as $N_{i}=\delta N_{i+1}$, for $11\leq i\leq m-$ 1, such that $\sum_{i=1}^{m}N_{i}\approx N$.
@@
  • Optimal progression. The evolution of samples follows the “optimal” geometric progression $N_{i}=2^{-(2k+d)i/2}N_{i+1}$ determined by Proposition 2.1.
Note that while Proposition 2.1 provides an optimal choice for the allocation of samples, the geometric factor intrinsically depends on the sampling cost assumption, which is architecture-dependent. We will compare the different sampling strategies described above in different numerical experiments in Section 3. The batch sizes $B_{m}\leq\ldots\leq B_{1}$ are defined through a geometric progression of the form $B_{i}=\delta B_{i+1}=\delta^{m-i}B_{m}$, where the batch size $B_{m}$ at the last level and the multiplier $\delta>1$ are prescribed manually.
Sampling strategy.
@@
  • Random sampling. We uniformly sample $\smash{\mathcal{B}_{2}^{k},\ldots\mathcal{B}_{m}^{k}}$ from the datasets $\smash{\mathcal{N}_{2},\ldots,\mathcal{N}_{m}}$. This is a standard multi-level Monte Carlo sampling and ensures independence of the data samples across levels in the telescopic sum given by Eq. 3.
@@
Fig. 6 in Section A.1 illustrates the two different sampling methods, where the finer data indices are uniformly samples from the original dataset (left), or sub-sampled from the previous resolution indices with the nested sub-sampling technique (right).
Multi-resolution data batcher.
@@ 2.3 MLMC optimization algorithm
Our MLMC optimization algorithm is described in Algorithm 1 and implements the mini-batch strategies introduced in Section 2.2, along with the telescopic sum evaluation of the gradient given by Eq. 3. At a given gradient correction level $2\leq i\leq m$, the loss is evaluated on samples $(a_{i,j},u_{i,j})$ from the batch $\mathcal{B}_{i}^{k}$ of the observation set $\mathcal{N}_{i}$ at resolution $R_{i}$. We then subtract the loss value at the same samples but coarsened to resolution $R_{i-1}$. The set of corresponding coarsen samples is denoted by $\tilde{\mathcal{N}}_{i-1}$. Note that it is essential to maintain consistency between the samples at different resolutions within a given correction level $i$ when constructing the data batcher, in order to preserve the validity of the MLMC estimator. Once the gradient has been computed, any gradient-based optimization algorithm can be used to update the parameters of the neural operator, such as Adam (Kingma and Ba, 2015).
@@ 3 Numerical experiments
We evaluate the efficiency MLMC training algorithm for neural operators on four state-of-the-art architectures: (i) Fourier neural operator (FNO) (Li et al., 2021b), (ii) message passing neural PDE solver (MP-PDE) (Brandstetter et al., 2022), and (iii) geometry-informed neural operator transformer (GINOT) (Liu et al., 2025). Additional experiments with (iv) mesh-free convolutional neural network (Zakariaei et al., 2025) are presented in Section C.1. The test-cases considered in this work consist of standard (time-dependent) nonlinear PDE problems which are summarized in Appendix B, and include a fluid flow past a cylinder, and a complex 3D jet engine brackets dataset (Hong et al., 2024). Note that the aim of this work is not to introduce new architectures, but rather improve the computational cost/accuracy trade-off in operator learning through MLMC training. All experiments are repeated on three random seeds and the results are averaged. The simulations are performed on a single Nvidia A100 GPU with 80GB of RAM.
@@ 3.1 Fourier neural operator
FNOs are a family of neural operators that exploit spectral (Fourier) discretizations of the input and output functions (Li et al., 2021b). These neural operators accept data at different resolutions using projection onto a latent dimension of fewer Fourier modes, which are then processed by Fourier convolutions. The MLMC training procedure introduced in Section 2 decreases the computational cost of the initial and final spectral convolutional layers and allows us to use fewer high-resolution data. This may be critical in large-scale applications where the cost of data acquisition might be high.
@@
We begin by training a two-dimensional FNO to learn the nonlinear diffusion coefficient field to solution operator associated with the standard Darcy flow problem introduced in (Li et al., 2021b). To ensure fairness of the experiments, we keep the same hyper-parameters employed in the code released by (Li et al., 2021b), and train the original FNO at different resolutions, which we refer to as the baseline below. We report in Fig. 1 the final test accuracy against average (over 500 epochs) computational training time per epoch of the standard FNO baseline as the data resolution increases from $15\times 15$ to $\{30\times 30,60\times 60,120\times 120,241\times 241\}$. We compare the results with the “Pareto curve” for different correction levels $m\in\{2,3,4,5\}$, resulting from MLMC training with decreasing sampling multiplier factor $\delta\in\{8,4,2,1.5,1\}$. As $\delta$ decreases, a higher number of fine-resolution data are used, which increases the accuracy but also the computational time. The proposed MLMC training algorithm enables us to reach almost the same testing accuracy achieved by traditional training at $R=241$, while requiring only $60\%$ of the training time. For completeness, we also report the tests referred to as “skip” in which the resolution scaling factor is $4$ (instead of $2$), and the considered resolutions are $15\times 15$, $60\times 60$, and $241\times 241$, as well as the nested sub-sampling strategy in Fig. 1 (right). In this way, we further increase the computational performance by reducing the total amount of fine samples required. Additional details about the dataset and numerical results are available in Sections B.1 and C.2.
@@
Figure 2: Navier-Stokes problem. Same as Fig. 1 with a FNO-3d trained on the time-dependent Navier-Stokes equations at increasing mesh resolutions $8\times 8$, $16\times 16$, $32\times 32$, $64\times 64$ (baseline) or with MLMC. The trained models are evaluated on a distinct test set at the finest resolution $64\times 64$. The magenta curve corresponds to MLMC at finest level $32\times 32$ and coarsest level $16\times 16$ with decreasing $\delta$, and yields a better cost/accuracy trade-off compared to the baseline.
Time-dependent Navier-Stokes equations.
@@
We then consider the approximation of the time-forward solution operator associated to the Navier-Stokes equations (see Section B.3) using a 3D Fourier neural operator (FNO-3d) that directly convolves in space-time. Also here, we employ the same hyper-parameters employed in the code released by (Li et al., 2021b). In Fig. 2, we display the Pareto frontier, in terms of average epoch training time and final testing loss, obtained by considering the standard FNO-3d training at different mesh resolutions (the baseline), and MLMC training with the two sampling strategies considered. Focusing on the baselines, we note that the testing loss values obtained at training resolutions $32\times 32$ and $64\times 64$ are very similar. Following this observation, we apply MLMC training with $m=2$ levels at finest resolution of $32\times 32$ and coarsest resolution of $16\times 16$, while decreasing the sampling factor $\delta$ from $8,4,2$. The result is illustrated as a magenta curve in Fig. 2, and yields the best trade-off in terms of accuracy versus computational training time. Surprisingly, it allows us to maintain the accuracy reached by the baseline at the resolution $R=64$, while decreasing the training time by a factor of $11$ (resp. $3$) compared to the baseline trained on a $64\times 64$ (resp. $32\times 32$) grid. We also considered the optimal sample allocation $N_{i}^{*}=4^{3i/2}N_{3}$ for $i\in\{1,2\}$ (predicted by the theory) with $\delta^{*}=2$, $N_{3}^{*}=45$ at resolution $32\times 32$, and plotted as a grey dot in Fig. 2. Finally, we do not observe a significant difference in terms of accuracy when considering nested sub-sampling over random sampling. Now considering Fig. 3(left), we compare ground truth solutions with the approximations obtained by MLMC training algorithm with $m=4$ (resp. $m=2$) levels and $\delta=4$ (resp. $\delta=2$) at the time $T=40$. Fig. 3 displays the evolution of the coarse and pairwise difference training losses composing the telescopic sum in Eq. 2 before evaluating the gradients for $m=4$ levels and $\delta=2$. As expected, the coarse loss is dominating the telescopic sum throughout the training, while the pairwise difference losses across levels decreases with the level. We note a very small contribution from the last level, explaining the improved performance of the magenta curve settings (maximum resolution of $32\times 32$) in Fig. 2.

### v2

@@ 2.2 Mini-batch strategies
$\mathbb{E}_{a\sim\mu}[\nabla_{\theta}X_{\theta}^{m}(a)]\approx\underbrace{\frac{1}{B_{1}}\sum_{j=1}^{B_{1}}[\nabla_{\theta}X_{\theta}^{1}(a_{1,j})]}_{\textnormal{coarsening loss}}+\sum_{i=2}^{m}\underbrace{\left(\frac{1}{B_{i}}\sum_{j=1}^{B_{i}}[\nabla_{\theta}X_{\theta}^{i}(a_{i,j})-\nabla_{\theta}X_{\theta}^{i-1}(a_{i,j})]\right)}_{\textnormal{pairing losses}},$ (6)
where $B_{i}$ stands for the batch size at the correction level $i$ in the sum hierarchy, obtained by drawing samples from a subset of $N_{i}$ and $N_{i-1}$ observations at resolution $R_{i}$ and $R_{i-1}$, respectively, for $2\leq i\leq m$. The term $\nabla_{\theta}X_{\theta}^{i}(a_{i,j})-\nabla_{\theta}X_{\theta}^{i-1}(a_{i,j})$ in Eq. (6) is therefore the gradient correction at level $i$, i.e., the difference between gradients evaluated on two consecutive discretization levels $i$ and i1i-1, for $2\leq i\leq m$. For clarity, throughout the paper we use level $i$ to indicate both the mesh resolution index and its associated gradient correction in the MLMC estimator. Moreover, the term $\nabla_{\theta}X_{\theta}^{i-1}(a_{i,j})$ in Eq. 6 is computed by coarsening the sample $a_{i,j}$ to resolution $R_{i-1}$ and evaluating the neural operator. Following Section 2.1, we decrease $B_{i}$ and $N_{i}$ with respect to the level $i$ as the data resolution $R_{i}$ increases.
Sample and batch sizes
@@
  • Geometric progression. There is a constant growing factor $\delta\geq 1$ between a level $i$ and a subsequent level $i+1$ as $N_{i}=\delta N_{i+1}$, for $1\leq i\leq m-1$, such that $\sum_{i=1}^{m}N_{i}\approx N$.
@@
  • Optimal progression. The evolution of samples follows the “optimal” geometric progression $N_{i}=2^{-(2k+d)i/2}N_{i-1}$, for $i=2,\ldots,m$, determined by Proposition 2.2.
Note that while Proposition 2.2 provides an optimal choice for the allocation of samples, the geometric factor intrinsically depends on the sampling cost assumption, which is architecture-dependent. We will compare the different sampling strategies described above in different numerical experiments in Section 3. The batch sizes $B_{m}\leq\ldots\leq B_{1}$ are defined through a geometric progression of the form $B_{i}=\delta B_{i+1}=\delta^{m-i}B_{m}$, where the batch size $B_{m}$ at the last level and the multiplier $\delta\geq 1$ are prescribed manually.
Sampling strategy
@@
  • Random sampling. We uniformly sample $\smash{\mathcal{B}_{2}^{k},\ldots\mathcal{B}_{m}^{k}}$ from the datasets $\smash{\mathcal{N}_{2},\ldots,\mathcal{N}_{m}}$. This is a standard multi-level Monte Carlo sampling and ensures independence of the data samples across levels in the telescopic sum given by Eq. 6. We repeat the sampling procedure for every epoch to ensure full coverage of the data.
@@
In Figs. 2 and 3 we report an illustration of the random sampling and nested sub-sampling strategies. In particular, we consider $N=105$ samples, $m=3$ levels ($R_{1}=4\times 4,R_{2}=8\times 8,R_{3}=16\times 16$), a multiplier $\delta=2$ and the geometrical progression. Data samples are selected at each level $i$, i.e. resolution, by obtaining the sets $\mathcal{N}_{1},\mathcal{N}_{2},\mathcal{N}_{3}$ of dimension $N_{1}=60,N_{2}=30,N_{3}=15$ and batches $\mathcal{B}_{1}^{k},\mathcal{B}_{2}^{k},\mathcal{B}_{3}^{k}$, $k=1,\ldots,3$, of dimension $B_{1}=20,B_{2}=10,B_{3}=5$.
Figure 2: Illustration of random sampling strategy in the construction of $\mathcal{B}$ for $m=3$ levels.
Fig. 2 illustrates the random sampling strategy, where the finer data indices are uniformly sampled from the original dataset. Note that for each resolution, it is possible to select samples that are not considered in the batches referring to a different resolution. The sample count refers to the number of times a sample is present in $\mathcal{B}$, not only because it can be sampled in another batch with the same resolution, but also because it can be chosen as a sample to be used in a correction term. To give some intuition, we start by uniformly sampling the indices of data related to $R_{1}=4\times 4$ and belonging to $\mathcal{B}^{k}_{1}$, with $k=1,\ldots,3$, each of dimension $B_{1}=20$. Then, we proceed be selecting the indices for the gradient correction at level $i=2$, that is batches $\mathcal{B}^{k}_{2}$, with $k=1,\ldots,3$, with batch size $B_{2}=10$. The same indices must also be selected at resolution $R_{1}$ in order to be able to compute the gradient correction term. The mechanism in then iterated to compute the gradient correction term at level $i=3$.
Figure 3: Illustration of nested sub-sampling strategy in the construction of $\mathcal{B}$ for $m=3$ levels.
In Fig. 3 finer data are sub-sampled from the previous resolution indices with the nested sub-sampling technique. In Fig. 3 (left), we show the selections of data indices at different resolutions; it is not possible to have a sample at level $i=3$ that has not also been selected at previous levels $i=1,2$. After obtaining the selected data indices, we perform a rebasing operation to better show the structure of batches $\mathcal{B}_{1}^{k},\mathcal{B}_{2}^{k},\mathcal{B}_{3}^{k}$, $k=1,\ldots,3$, resulting from nested sub-sampling in Fig. 3 (right).
Multi-resolution data batcher
@@ 2.3 MLMC optimization algorithm
Our MLMC optimization algorithm is described in Algorithm 1 and implements the mini-batch strategies introduced in Section 2.2, along with the telescopic sum evaluation of the gradient given by Eq. 6. At a given gradient correction level $2\leq i\leq m$, the loss is evaluated on samples $(a_{i,j},u_{i,j})$ from the batch $\mathcal{B}_{i}^{k}$ of the observation set $\mathcal{N}_{i}$ at resolution $R_{i}$. We then subtract the loss value at the same samples but coarsened to resolution $R_{i-1}$. The set of corresponding coarsen samples is denoted by $\tilde{\mathcal{N}}_{i-1}$. Note that it is essential to maintain consistency between the samples at different resolutions within a given correction level $i$ when constructing the data batcher, in order to preserve the validity of the MLMC estimator and compute the correct pairing losses in Eq. 6. Once the gradient has been computed, any gradient-based optimization algorithm can be used to update the parameters of the neural operator, such as Adam [21].
@@ 3 Numerical experiments
We evaluate the efficiency of the MLMC training algorithm for neural operators on three state-of-the-art neural operators architectures: (i) Fourier neural operator (FNO) [32], (ii) message passing neural PDE solver (MP-PDE) [5], and (iii) geometry-informed neural operator transformer (GINOT) [33]. Additional experiments with (iv) mesh-free convolutional neural network [44] are also reported. The test-cases considered in this work consist of standard (time-dependent) nonlinear PDE problems, and include a fluid flow past a cylinder, and a complex 3D jet engine brackets dataset [19]. Note that the aim of this work is not to introduce new architectures, but rather improve the computational cost/accuracy trade-off in operator learning through MLMC training. All experiments are repeated on three random seeds and the results are averaged. The simulations are performed on a single Nvidia A100 GPU with 80GB of RAM. Code to reproduce all experiments in this study is openly available at https://github.com/JRowbottomGit/mlmc_optim.
@@ 3.1 Fourier neural operator
FNOs are a family of neural operators that exploit spectral (Fourier) discretizations of the input and output functions [32]. These neural operators accept data at different resolutions using projection onto a latent dimension of fewer Fourier modes, which are then processed by Fourier convolutions. The MLMC training procedure introduced in Section 2 decreases the computational cost of the initial and final spectral convolutional layers and allows us to use fewer high-resolution data. This may be critical in large-scale applications where the cost of data acquisition might be high.
Darcy flow problem
We consider the steady state Darcy flow equation on the domain $\Omega=[0,1]^{2}$ as [32]:
$u(\mathbf{x})$ $=0,\quad\mathbf{x}\in\partial\Omega,$
where the diffusion coefficient $a:\Omega\rightarrow\mathbb{R}$ is chosen such that $a(\mathbf{x})=\psi\circ\mu(\mathbf{x})$, where $\mu\sim\mathcal{N}(0,(-\Delta+9I)^{-2})$ with zero Neumann boundary conditions on the Laplacian, and $\psi:\mathbb{R}\to\mathbb{R}$ is defined as $\psi(x)=12$ if $x\geq 0$ and $3$ otherwise. We are interested in learning the nonlinear operator $a\mapsto u$ mapping the diffusion operator $a$ to the corresponding solution $u$ to Eq. 7. The equation is solved numerically for $N=1024$ random diffusion coefficients on an equispaced $241\times 241$ grid using a second-order finite difference scheme, and downsampled to the lower resolutions such that $\{R_{i}\times R_{i}\}_{i=1}^{5}=\{15\times 15,30\times 30,60\times 60,120\times 120,241\times 241\}$.
@@
To ensure fairness of the experiments, we keep the same hyper-parameters employed in the code released by [32], and train the original FNO at different resolutions, which we refer to as the baseline below. We report in Fig. 4 the final test accuracy against average (over 500 epochs) computational training time per epoch of the standard FNO baseline as the data resolution increases from $15\times 15$ to $\{30\times 30,60\times 60,120\times 120,241\times 241\}$. We compare the results with the “Pareto curve” for different number of levels $m\in\{2,3,4,5\}$, resulting from MLMC training with decreasing sampling multiplier factor $\delta\in\{8,4,2,1.5,1\}$. As $\delta$ decreases, a higher number of fine-resolution data are used, which increases the accuracy but also the computational time. The proposed MLMC training algorithm enables us to reach almost the same testing accuracy achieved by traditional training at $R=241$, while requiring only $60\%$ of the training time. For completeness, we also report the tests referred to as “skip” in which the resolution scaling factor is $4$ (instead of $2$), and the considered resolutions are $15\times 15$, $60\times 60$, and $241\times 241$, as well as the nested sub-sampling strategy in Fig. 4 (right). In this way, we further increase the computational performance by reducing the total amount of fine samples required.
[…]
Moreover, we test the generalization capabilities of the FNO, trained by means of the MLMC algorithm, on the finest resolution $R=241$, by varying the training budget. This analysis is motivated by the interest to evaluate the ability of the MLMC training algorithm to improve resolution generalization. We considered $m=2$ levels, resolutions $\{R_{i}\times R_{i}\}_{i=1}^{2}=\{30\times 30$, $120\times 120\}$, resampling multiplier factor $\delta=2,4$ and the nested sub-sampling technique to test the generalization performance on the finest resolution $R=241$. The results are reported in Fig. 6 where we report the final testing loss against the cumulative training computational time for the baseline with $R=120$ and the MLMC training algorithm approximation. The MLMC training algorithm enables to improve resolution generalization by achieving a prescribed accuracy for a lower computational budget.
Figure 6: Darcy problem. Final testing loss against cumulative training computational time.
Finally, we compare the training computational cost/accuracy trade-off enabled by the MLMC training algorithm employing the gradient estimation formula in Eq. 3 with respect to the one obtained by means of a mixed-resolution batches training without telescoping corrections. In particular, we perform the test by using the training specifics $m=4$ and $\delta=2$. The average training computational time obtained by the mixed-resolution batches approach is $1.13s$ with a final testing loss equal to $0.022$. In contrast, with the MLMC training algorithm we obtain the same accuracy but requireing just the $37\%$ of training computational time.
Time-dependent Navier-Stokes equations
We then consider the approximation of the time-forward solution operator associated to the Navier-Stokes equations using a 3D Fourier neural operator (FNO-3d) that directly convolves in space-time. The two-dimensional time-dependent Navier-Stokes equations are expressed in vorticity form on the unit torus as follows [32]:
$w(\mathbf{x},0)$ $=w_{0}(\mathbf{x}),$ $\mathbf{x}\in(0,1)^{2},$
where $\mathbf{u}:\Omega\times[0,T)\rightarrow\mathbb{R}^{2}$ is the velocity field and $w=\nabla\times\mathbf{u}$ is the vorticity field. The initial condition for the vorticity field is generated as $w_{0}\sim\mathcal{N}\bigl(0,\,7^{3/2}(-\Delta+49I)^{-2.5}\bigr)$, the forcing term is given by $f=0.1(\sin(2\pi(x+y))+\cos(2\pi(x+y)))$, and the viscosity coefficient is equal to $\nu=10^{-3}$. Eq. 8 are equipped with periodic boundary conditions. We are interested in learning the operator which maps the vorticity field up to time $t=10$ to the vorticity up to time $T=50$. The equations are numerically solved by a pseudo-spectral method and a Crank–Nicolson time scheme for the viscous term with an explicit treatment of the nonlinear term on a uniform grid of resolution $256\times 256$, and a time discretization step $\Delta t=10^{-4}$. The dataset is generated by downsampling the solutions to $\{R_{i}\times R_{i}\}_{i=1}^{4}=\{8\times 8$, $16\times 16$, $32\times 32,64\times 64\}$ space discretization points and storing them at every time units $t\in\{1,\ldots,50\}$. The number of samples of the initial conditions $\omega_{0}$ and resulting trajectories is set to $N=3000$.
@@
Figure 7: Navier-Stokes problem. Same as Fig. 4 with an FNO-3d trained on the time-dependent Navier-Stokes equations at increasing mesh resolutions $8\times 8$, $16\times 16$, $32\times 32$, $64\times 64$ (baseline) or with MLMC. The trained models are evaluated on a distinct test set at the finest resolution $64\times 64$. The magenta curve corresponds to MLMC at finest level $32\times 32$ and coarsest level $16\times 16$ with decreasing $\delta$, and yields a better cost/accuracy trade-off compared to the baseline.
@@
Again, we employ the same hyper-parameters employed in the code released by [32]. In Fig. 7, we display the Pareto frontier, in terms of average epoch training time and final testing loss, obtained by considering the standard FNO-3d training at different mesh resolutions (the baseline), and MLMC training with the two sampling strategies considered. Focusing on the baselines, we note that the testing loss values obtained at training resolutions $32\times 32$ and $64\times 64$ are very similar. Following this observation, we apply MLMC training with $m=2$ levels at finest resolution of $32\times 32$ and coarsest resolution of $16\times 16$, while decreasing the sampling factor $\delta$ from $8,4,2$. The result is illustrated as a cyan curve in Fig. 7, and yields the best trade-off in terms of accuracy versus computational training time. Surprisingly, it allows us to maintain the accuracy reached by the baseline at the resolution $R=64$, while decreasing the training time by a factor of $11$ (resp. $3$) compared to the baseline trained on a $64\times 64$ (resp. $32\times 32$) grid. We also considered the optimal sample allocation $N_{i}^{*}=2^{2i}N_{3}$ for $i\in\{1,2\}$ with $\delta^{*}=2$, $N_{3}^{*}=45$ at resolution $32\times 32$, and plotted as a bordeaux dot in Fig. 7. Finally, we do not observe a significant difference in terms of accuracy when considering nested sub-sampling over random sampling. Now considering Fig. 8 (left), we compare ground truth solutions with the approximations obtained by MLMC training algorithm with $m=4$ (resp. $m=2$) levels and $\delta=4$ (resp. $\delta=2$) at the time $T=40$. Fig. 8 displays the evolution of the coarse and pairwise difference training losses composing the telescopic sum in Eq. 3 before evaluating the gradients for $m=4$ levels and $\delta=2$. As expected, the coarse loss is dominating the telescopic sum throughout the training, while the pairwise difference losses across levels decreases with the level. We note a very small contribution from the last level, explaining the improved performance of the magenta curve settings (maximum resolution of $32\times 32$) in Fig. 7.